OpenStudy (anonymous):

HELP PLEASE!!!!!!!!!!!! Solve the following quadratic equation by factoring:

OpenStudy (anonymous):

\[x ^{2}-10x-24=0\]

OpenStudy (aravindg):

try to find two numbers whose sum is -10 and product is 24

OpenStudy (aravindg):

@nburgher

OpenStudy (anonymous):

2*-12

OpenStudy (anonymous):

@AravindG

OpenStudy (aravindg):

so we can write the given equation as (x+2)(x-12)=0

OpenStudy (aravindg):

and hence factorised :)

OpenStudy (anonymous):

yes x=2 or x=-12 but when put into a graphing calculator it does not make a parobola graph

OpenStudy (anonymous):

when doing quadratic equations does it always have to be a parobola graph @AravindG

OpenStudy (aravindg):

it will !!

OpenStudy (aravindg):

yes a quadratic will always be a prabola

OpenStudy (anonymous):

it does not I checked it on a graphing calculator

OpenStudy (anonymous):

it makes a straight line

OpenStudy (aravindg):

try like this \[(x+2) \times (x-12)=0\]

OpenStudy (aravindg):

try that

OpenStudy (aravindg):

multiplication should be there

OpenStudy (anonymous):

it is still a straight line and it shows the x-intercept as -2

OpenStudy (anonymous):

@AravindG

OpenStudy (anonymous):

(-2,0)

OpenStudy (anonymous):

never mind I had to zoom out on the calculator

OpenStudy (aravindg):

my graphing calculator is THE BEST :p http://www.wolframalpha.com/input/?i=%28x%2B2%29%28x-12%29%3D0

OpenStudy (anonymous):

thank you for your help. I gave you a medal

OpenStudy (aravindg):

yw :)